OpenStudy (anonymous):

I'LL GIVE A MEDAL FOR THE BEST ANSWER A rhombus shaped mirror has diagonals of 36 inches & 48 inches. What is the area of this mirror?

OpenStudy (anonymous):

@jdoe0001

jimthompson5910 (jim_thompson5910):

Use the formula A = (p*q)/2 where A = area p = length of one diagonal q = length of the other diagonal

OpenStudy (calculusxy):

Area=b*h

OpenStudy (calculusxy):

I believe same as rectangle and square.

OpenStudy (anonymous):

thanks @calculusxy & @jim_thompson5910